OpenStudy (anonymous):

The point (−0.5,−2) is the midpoint of (−5,−8) and what point?

OpenStudy (mertsj):

\[\frac{x-5}{2}=-.5; \frac{y-8}{2}=-2\]

OpenStudy (anonymous):

CAN YOU TELL ME THE MISSING POINT?

OpenStudy (mertsj):

If you solve the two equations I wrote you will have answers for x and y and that will be the point you desire.

OpenStudy (anonymous):

(4,4)

OpenStudy (anonymous):

I am not confident in my answer, (4,4)

OpenStudy (anonymous):

The equation of a circle C is x2+y2−18x−2y+57=0. What is its center (h,k) and its radius r?